hiiiihiiiiiiiiii:

Divide: 2 1/3÷1/9. Reduce to lowest terms. None of the answers is correct. 21 63/3 7/27

MxxnLight:

Maybe @Vocaloid can help you? or @Laylalyssa

SemiDefinite:

2 1/3÷1/9 = 7/3 ÷ 1/9 = 7/3 * 9/1